OpenStudy (anonymous):

Which equation has no solution? a. 4(x + 3) + 2x = 6(x + 2) b. 5 + 2(3 + 2x) = x + 3(x + 1) c. 5(x + 3) + x = 4(x + 3) + 3 d. 4 + 6(2 + x) = 2(3x + 8)

OpenStudy (anonymous):

the way i'd do it is to go one by one putting it in a mx+b form lemme do a for you

OpenStudy (anonymous):

4(x + 3) + 2x = 6(x + 2) 4x+12+2x=6x+12 distribute the 4 on the left hand side then the 6 on RHS 6x+12=6x+12 add like terms now since both sides are equal i think they have infinitely many solutions. SO you move on

OpenStudy (anonymous):

to summarize 1. You distribute the number outside the parenthesis in both sides of the equation. 2. add the like terms in both sides of the equation 3. you will end up with something looking like mx+b on both sides there are 3 possibilities. and keep in mind x is a variable mx+b=mx+b same m and same b : infinite solutions mx+b=Mx+b different m same b : one solution mx+b =mx+B different b same m : no solution ANSWER

OpenStudy (anonymous):

So the answer is

OpenStudy (anonymous):

??

OpenStudy (anonymous):

APply what i said to B C D

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

you can do it just try B :)

OpenStudy (anonymous):

Whats the answer

OpenStudy (anonymous):

you wanna know how to do it or just want the answer?

OpenStudy (anonymous):

the answer please

OpenStudy (anonymous):

@dronchik

OpenStudy (mathmath333):

5 + 2(3 + 2x) = x + 3(x + 1) \(\Large\tt \color{black}{4x+11=4x+3}\) here if u substitute x=1 then \(\Large\tt \color{black}{4+11=4+3}\) \(\Large\tt \color{black}{15\neq7}\) or if u substitute x=2 \(\Large\tt \color{black}{8+11=8+3}\) \(\Large\tt \color{black}{19\neq11}\) \(\tt \color{black}{\text{so due to contradiction this equation has no solution }}\)